OpenStudy (anonymous):

Which of the following equations could be the result of using the comparison method to solve the system shown? x + y = 5 2x + y = 7 -x - 5 = 7 - 2x 5 - x = 7 - 2x 5 - x = 2x - 7 got b

Directrix (directrix):

I don't recognize the term " comparison method" but may know it as Addition/Subtraction or something else. What is " comparison method?" Let me know and then I'll see what I get as an answer.

OpenStudy (anonymous):

one sec.... let me find it in my lesson

OpenStudy (anonymous):

Example: Solve the following system of equations by the comparison method. 2x + 6y + 3 = 0 x - 4y - 9 = 0 Solution: Make y the subject of each equation; then solve each equation: Reminder: Cross-multiply to solve a proportion. Combine the two results into a single equation of the results and find the value of x: Substitute 3 in for x into one of the original equations and solve for y: 2(3) + 6y + 3 = 0 6 + 6y + 3 = 0 6y + 9 = 0 6y = -9 y = - The checks were shown above. Therefore, the solution set is {(3, -)}

OpenStudy (anonymous):

crap the fractions didnt show up

Directrix (directrix):

x + y = 5 --> y = 5 -x 2x + y = 7 --> y = 7 - 2x -------------------------- So, 5-x = 7 - 2x ---------------------> This is what I got. That technique is what I know as substitution. That may not be the same as comparison.

Directrix (directrix):

According to the instructions and the options given, the answer is not the solution to the two equations but just what you get after you set the two y expressions equal. >>>Solution: Make y the subject of each equation; So, the answer I see as correct is the one you asked about, B. 5 - x = 7 - 2x